$SPAD/src/input coordsys.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- test input for CoordinateSystems package
draw(sin(x),x=0.5..%pi,coordinates == bipolar(1$DFLOAT))
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,0..2*%pi, 0..%pi,coordinates == bipolar(1$DFLOAT))
draw(surface(u*cos(v),u*sin(v),u),u=1..4,v=1..2*%pi,coordinates == _
 bipolarCylindrical(1$DFLOAT))
--conical(a,b) maps a 3D point (lambda,mu,nu) to the rectangular coordinates:
--x = lambda*mu*nu/(a*b)
--y = lambda/a*sqrt((mu**2-a**2)*(nu**2-a**2)/(a**2-b**2))
--z = lambda/b*sqrt((mu**2-b**2)*(nu**2-b**2)/(b**2-a**2))
--NOTE:  There will be a division by zero error if a*b = 0, or a**2-b**2 = 0,
--     or if b**2-a**2 = 0. Also, the following relations must be true:
--       (mu**2-a**2)*(nu**2-a**2)/(a**2-b**2) > 0 and
--       (mu**2-b**2)*(nu**2-b**2)/(b**2-a**2) > 0.
j1(t:DFLOAT):DFLOAT == 4
j2(t:DFLOAT):DFLOAT == t
draw(curve(j1,j2,j2),-9..9,coordinates == cylindrical)
draw(sin(4*t/7),t=0..14*%pi,coordinates == elliptic(1$DFLOAT))
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,0..2*%pi,0..%pi,coordinates == elliptic(1$DFLOAT))
U2:Vector Expression Integer := vector [0,0,1]
x(u,v) == beta(u) + v*delta(u)
beta u == vector [cos u, sin u, 0]
delta u == (cos(u/2)) * beta(u) + sin(u/2) * U2
vec := x(u,v)
draw(surface(vec.1,vec.2,vec.3),v=-0.5..0.5,u=0..2*%pi,coordinates == _
  ellipticCylindrical(1$DFLOAT),var1Steps == 50,var2Steps == 50)
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,-%pi/2..%pi/2,0..2*%pi,coordinates == oblateSpheroidal(1$DFLOAT))
h1(t:DFLOAT):DFLOAT == t
h2(t:DFLOAT):DFLOAT == 2
draw(curve(h1,h2),-3..3,coordinates == parabolic)

draw(surface(u*cos(v),u*sin(v),2*u),u=0..4,v=0..2*%pi,coordinates == _
  parabolic)
draw(surface(u*cos(v),u*sin(v),v*cos(u)),u=0..4,v=0..2*%pi,coordinates == _
  parabolicCylindrical)
draw(surface(u*cos(v),u*sin(v),u*v),u=0..4,v=0..2*%pi,coordinates == _
  paraboloidal,var1Steps == 50, var2Steps == 50)
draw(sin(4*t/7),t=0..14*%pi,coordinates == polar)
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,0..2*%pi, 0..%pi,coordinates == polar)
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,-%pi/2..%pi/2,0..2*%pi,coordinates == prolateSpheroidal(1$DFLOAT))
m(u:DFLOAT,v:DFLOAT):DFLOAT == 1
draw(m,0..2*%pi,0..%pi,coordinates == spherical)
draw(surface(u*cos(v),u*sin(v),u),u=1..4,v=1..4*%pi,coordinates == _
  toroidal(1$DFLOAT))

———-

References^{[1] nothing}

5